# Efficient Perplexity Bound and Ratio Matching in Discrete Diffusion Language Models

### v1

@@ A.1 KL Divergence Theorems
Let $q^{*}_{1}$ and $\hat{q}_{1}$ denote two initial distributions and define two Continuous-Time Markov Chains (CTMCs) running from time $1$ to 00, with transition-rate matrices ${\bm{Q}}^{*}_{t}(y,x)={\bm{Q}}_{t}(x,y)r(x,y)$ and $\hat{{\bm{Q}}}_{t}(y,x)={\bm{Q}}_{t}(x,y)\hat{r}(x,y)$, where $r(x,y)$ and $\hat{r}(x,y)$ are chosen such that they ensure ${\bm{Q}}^{*}_{t}$ and $\hat{{\bm{Q}}}_{t}$ are indeed transition-rate matrices. When the first process (with matrix ${\bm{Q}}^{*}_{t}(y,x)$) is applied to initial distribution $q^{*}_{1}$ and the second (with matrix $\hat{{\bm{Q}}}_{t}(y,x)$) to initial distribution $\hat{q}_{1}$, they define distributions $q_{t}^{*}$ and $\hat{q}_{t}$ at time $t$, and $q_{0}^{*}$, $\hat{q}_{0}$ at the end of the diffusion, for which
@@
From the data processing inequality applied to the reverse of the processes defined in the Lemma, we have $D_{KL}(q_{0}^{*}||\hat{q}_{0})\leq D_{KL}(q^{*}||\hat{q})$ where $q^{*}$ is the measure over the space of paths generated by the first process, while $\hat{q}$ is the measure over the space of paths generated by second one. Using the expression for the KL divergence between path measures of two CTMCs found in (Opper & Sanguinetti, 2007, Section 2.1), and by substituting $q_{t}(x):=q^{*}(x_{t})$, $g(x^{\prime}|x)={\bm{Q}}^{*}_{t}(y,x_{t})$ and $f(x^{\prime}|x)=\hat{{\bm{Q}}}_{t}(y,x_{t})$, we have:
@@
Theorem 1. Define a CTMC with transition matrix ${{\bm{Q}}}_{t}$ that runs from time 00 to $1$. The true reverse process defines a probability evolution $p_{t}$ from $p_{1}$ to the data distribution $p_{0}$, while the learned reverse process induces the evolution $p^{\theta}_{t}$ from the reference distribution $p^{\theta}_{1}=p_{r}$ to the approximation of the data distribution $p^{\theta}_{0}$. In this setting, the following KL divergence bound holds
@@
In Lemma 1, we set $r(x,y)=\frac{p_{t}(y)}{p_{t}(x)}$ and $\hat{r}(x_{t},y)=s_{\theta}(x_{t},t)_{y}$. This implies that $q_{t}^{*}=p_{t}$ and $\hat{q}_{t}=p_{t}^{\theta}$ In particular, $q_{0}^{*}=p_{0}$, $\hat{q}_{0}=p_{0}^{\theta}$ and $q_{1}^{*}=p_{1}$, $\hat{q}_{1}=p_{1}^{\theta}:=p_{r}$. Plugging everything in Expression (20) gives the desired result. ∎
@@
$-\log p_{0}^{\theta}(x_{0})\leq\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot|x_{0})}\sum_{y\neq x_{t}}{\bm{Q}}_{t}(x_{t},y)\ell\left(\frac{p_{t|0}(y|x_{0})}{p_{t|0}(x_{t}|x_{0})},s_{\theta}(x_{t},t)_{y}\right)dt$
$+D_{KL}(p_{1|0}(\cdot|x_{0})||p_{r}).$ (28)
@@
In Lemma 1 we set $r(x,y)=\frac{p_{t|0}(y|x_{0})}{p_{t|0}(x_{t}|x_{0})}$ and $\hat{r}(x_{t},y)=s_{\theta}(x_{t},t)_{y}$, where $p_{t|0}$ is the probability over states at time $t$ determined by a CTMC with transition-rate matrix ${\bm{Q}}_{t}$ applied to initial distribution $\delta(x=x_{0})$. As such, we have $q_{t}^{*}=p_{t|0}$ and $\hat{q}_{t}=p_{t}^{\theta}$. In particular, $q_{0}^{*}=\delta(x=x_{0})$, $\hat{q}_{0}=p_{0}^{\theta}$ and $q_{1}^{*}=p_{1|0}$ , $\hat{q}_{1}=p_{1}^{\theta}:=p_{r}$. Plugging everything in Expression (20):
$D_{KL}(\delta(x=x_{0})||p_{0}^{\theta})\leq\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot|x_{0})}\sum_{y\neq x_{t}}{\bm{Q}}_{t}(x_{t},y)\left[K\left(\frac{p_{t|0}(y|x_{0})}{p_{t|0}(x_{t}|x_{0})}\right)\right.\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad$
@@
[…]
@@
[…]
@@
$=-\left(\int\sum_{y}{\bm{Q}}_{t}(x_{t},y)p_{t}(y)\log(p^{\theta}_{t}(x_{t}))dx_{t}+\int\frac{p_{t}(x_{t})}{p^{\theta}_{t}(x_{t})}\sum_{y}{\bm{Q}}_{t}(x_{t},y)p^{\theta}_{t}(y)dx_{t}\right)$ (33)
$=-\left(\int\sum_{y}{\bm{Q}}_{t}(x_{t},y)\frac{p_{t}(y)}{p_{t}(x_{t})}\log(p^{\theta}_{t}(x_{t}))p(x_{t})dx_{t}+\int p_{t}(x_{t})\sum_{y}{\bm{Q}}_{t}(x_{t},y)\frac{p^{\theta}_{t}(y)}{p^{\theta}_{t}(x_{t})}dx_{t}\right)$ (34)
$=-\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)\left(\frac{p_{t}(y)}{p_{t}(x_{t})}\log(p^{\theta}_{t}(x_{t}))+\frac{p^{\theta}_{t}(y)}{p^{\theta}_{t}(x_{t})}\right)$ (35)
$=-\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)\left(\frac{p^{\theta}_{t}(y)}{p^{\theta}_{t}(x_{t})}-\frac{p_{t}(y)}{p_{t}(x_{t})}\log\frac{p^{\theta}_{t}(y)}{p^{\theta}_{t}(x_{t})}\right)$ (36)
$+\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)[\frac{p_{t}(y)}{p_{t}(x_{t})}\log p^{\theta}_{t}(y)].$ (37)
@@
$=\sum_{x_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)p_{t}(y)\log(p^{\theta}_{t}(y))=\sum_{y}\left(\sum_{x_{t}}{\bm{Q}}_{t}(x_{t},y)\right)p_{t}(y)\log(p^{\theta}_{t}(y))=0,$ (39)
@@
$=\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)\left(\frac{p^{\theta}_{t}(y)}{p^{\theta}_{t}(x_{t})}-\frac{p_{t}(y)}{p_{t}(x_{t})}\log\frac{p^{\theta}_{t}(y)}{p^{\theta}_{t}(x_{t})}\right)dt.$ (41)
@@
$-H(p_{0})+H(p_{1})=\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)K\left(\frac{p_{t}(y)}{p_{t}(x_{t})}\right)dt,$ (45)
@@
$=\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)\left[K\left(\frac{p_{t}(y)}{p_{t}(x_{t})}\right)+\left(s_{\theta}(x_{t},t)_{y}-\frac{p_{t}(y)}{p_{t}(x_{t})}\log{s_{\theta}(x_{t},t)_{y}}\right)\right]dt.$ (47)
@@
$=\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y\neq x_{t}}{\bm{Q}}_{t}(x_{t},y)\left[K\left(\frac{p_{t}(y)}{p_{t}(x_{t})}\right)+\left(s_{\theta}(x_{t},t)_{y}-\frac{p_{t}(y)}{p_{t}(x_{t})}\log{s_{\theta}(x_{t},t)_{y}}\right)\right]dt.$ (50)
@@
$-\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y\neq x_{t}}{\bm{Q}}_{t}(y,x_{t})dt+H(p_{1},p_{r}),$ (51)
where ¯$\bar{\ell}(a,b)=\left(b-a\log b\right)$.
@@
$+\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)\left(s_{\theta}(x_{t},t)_{y}-\frac{p_{t}(y)}{p_{t}(x_{t})}\log{s_{\theta}(x_{t},t)_{y}}\right)dt+D_{KL}(p_{1}||p_{r}).$ (55)
@@
[…]
@@
$-\int_{0}^{1}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}}\sum_{{\bm{y}}\neq{\bm{x}}_{t}}{\bm{Q}}_{t}({\bm{y}},{\bm{x}}_{t})dt=-\mathbb{E}_{t\sim U(0,1)}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\sum_{i=1}^{L}\sum_{{\bm{y}}^{i}\neq{\bm{x}}^{i}_{t}}{\bm{Q}}_{t}^{tok}({\bm{y}}^{i},{\bm{x}}^{i}_{t})$ (58)
$=-\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\sum_{{\bm{y}}^{i}\neq{\bm{x}}^{i}_{t}}{\bm{Q}}_{t}^{tok}({\bm{y}}^{i},{\bm{x}}^{i}_{t})$ (59)
$=-\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\sum_{{\bm{x}}^{i}_{t}}p_{t}({\bm{x}}^{i}_{t})\sum_{{\bm{y}}^{i}\neq{\bm{x}}^{i}_{t}}{\bm{Q}}_{t}^{tok}({\bm{y}}^{i},{\bm{x}}^{i}_{t})$ (60)
$=-\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\sum_{{\bm{x}}^{i}_{0}}p_{0}({\bm{x}}^{i}_{0})\sum_{{\bm{x}}^{i}_{t}}p_{t|0}({\bm{x}}^{i}_{t}|{\bm{x}}^{i}_{0})\sum_{{\bm{y}}^{i}\neq{\bm{x}}^{i}_{t}}{\bm{Q}}_{t}^{tok}({\bm{y}}^{i},{\bm{x}}^{i}_{t})$ (61)
@@
$-\int_{0}^{1}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}}\sum_{{\bm{y}}\neq{\bm{x}}_{t}}{\bm{Q}}_{t}({\bm{y}},{\bm{x}}_{t})=$ (62)
$=-\left(1-\frac{1-p_{m}}{n-1}\right)\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\sum_{{\bm{x}}^{i}_{0}}p_{0}({\bm{x}}^{i}_{0})\sum_{{\bm{x}}^{i}_{t}\neq n}p_{t|0}({\bm{x}}^{i}_{t}|{\bm{x}}^{i}_{0})\sigma_{t}^{{}^{\prime}}=$ (63)
$=-\left(1-\frac{1-p_{m}}{n-1}\right)\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\sum_{{\bm{x}}^{i}_{0}}p_{0}({\bm{x}}^{i}_{0})\sigma_{t}^{{}^{\prime}}\sum_{{\bm{x}}^{i}_{t}\neq n}p_{t|0}({\bm{x}}^{i}_{t}|{\bm{x}}^{i}_{0})$ (64)
$=-\left(1-\frac{1-p_{m}}{n-1}\right)\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\sum_{{\bm{x}}^{i}_{0}}p_{0}({\bm{x}}^{i}_{0})\sigma_{t}^{{}^{\prime}}\left(1-p({\bm{x}}^{i}_{t}=n|{\bm{x}}^{i}_{0})\right)$ (65)
[…]
$=\left(1-\frac{1-p_{m}}{n-1}\right)\frac{L}{p_{m}}(e^{-\sigma_{1}p_{m}}-e^{-\sigma_{0}p_{m}})$ (67)
@@
[…]

### v2

@@ A.1 KL Divergence Theorems
Let $q^{*}_{1}$ and $\hat{q}_{1}$ denote two initial distributions and define two Continuous-Time Markov Chains (CTMCs) running from time $1$ to 0, with transition-rate matrices ${\bm{Q}}^{*}_{t}(y,x)={\bm{Q}}_{t}(x,y)r(x,y)$ and $\hat{{\bm{Q}}}_{t}(y,x)={\bm{Q}}_{t}(x,y)\hat{r}(x,y)$, where $r(x,y)$ and $\hat{r}(x,y)$ are chosen such that they ensure ${\bm{Q}}^{*}_{t}$ and $\hat{{\bm{Q}}}_{t}$ are indeed transition-rate matrices. When the first process (with matrix ${\bm{Q}}^{*}_{t}(y,x)$) is applied to initial distribution $q^{*}_{1}$ and the second (with matrix $\hat{{\bm{Q}}}_{t}(y,x)$) to initial distribution $\hat{q}_{1}$, they define distributions $q_{t}^{*}$ and $\hat{q}_{t}$ at time $t$, and $q_{0}^{*}$, $\hat{q}_{0}$ at the end of the diffusion, for which
@@
[…]
@@
Theorem 1. Define a CTMC with transition matrix ${{\bm{Q}}}_{t}$ that runs from time 0 to $1$. The true reverse process defines a probability evolution $p_{t}$ from $p_{1}$ to the data distribution $p_{0}$, while the learned reverse process induces the evolution $p^{\theta}_{t}$ from the reference distribution $p^{\theta}_{1}=p_{r}$ to the approximation of the data distribution $p^{\theta}_{0}$. In this setting, the following KL divergence bound holds
@@
In Lemma 1, we set $r(x,y)=\frac{p_{t}(y)}{p_{t}(x)}$ and $\hat{r}(x_{t},y)=s_{\theta}(x_{t},t)_{y}$. This implies that $q_{t}^{*}=p_{t}$ and $\hat{q}_{t}=p_{t}^{\theta}$ . In particular, $q_{0}^{*}=p_{0}$, $\hat{q}_{0}=p_{0}^{\theta}$ and $q_{1}^{*}=p_{1}$, $\hat{q}_{1}=p_{1}^{\theta}:=p_{r}$. Plugging everything in Expression (20) gives the desired result. ∎
@@
$\log p_{0}^{\theta}(x_{0})\leq\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot|x_{0})}\sum_{y\neq x_{t}}{\bm{Q}}_{t}(x_{t},y)\ell\left(\frac{p_{t|0}(y|x_{0})}{p_{t|0}(x_{t}|x_{0})},s_{\theta}(x_{t},t)_{y}\right)dt$
$D_{KL}(p_{1|0}(\cdot|x_{0})||p_{r}).$ (28)
@@
In Lemma 1 we set $r(x,y)=\frac{p_{t|0}(y|x_{0})}{p_{t|0}(x_{t}|x_{0})}$ and $\hat{r}(x_{t},y)=s_{\theta}(x_{t},t)_{y}$, where $p_{t|0}$ is the probability over states at time $t$ determined by a CTMC with transition-rate matrix ${\bm{Q}}_{t}$ applied to initial distribution $\delta(x=x_{0})$.
As such, we have $q_{t}^{*}=p_{t|0}$ and $\hat{q}_{t}=p_{t}^{\theta}$. In particular, $q_{0}^{*}=\delta(x=x_{0})$, $\hat{q}_{0}=p_{0}^{\theta}$ and $q_{1}^{*}=p_{1|0}$ , $\hat{q}_{1}=p_{1}^{\theta}:=p_{r}$. Plugging everything in Expression (20):
$D_{KL}(\delta(x=x_{0})||p_{0}^{\theta})\leq\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot|x_{0})}\sum_{y\neq x_{t}}{\bm{Q}}_{t}(x_{t},y)\left[K\left(\frac{p_{t|0}(y|x_{0})}{p_{t|0}(x_{t}|x_{0})}\right)\right$$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad$
@@
$H(p_{0},p_{0}^{\theta})-H(p_{1},p_{1}^{\theta})=\int_{1}^{0}\frac{\partial}{\partial t}H(p_{t},p_{t}^{\theta})dt$ (30)
@@
$\left(\int\frac{\partial p_{t}(x_{t})}{\partial t}\log(p^{\theta}_{t}(x_{t}))dx_{t}+\int\frac{p_{t}(x_{t})}{p^{\theta}_{t}(x_{t})}\frac{\partial p^{\theta}_{t}(x_{t})}{\partial t}dx_{t}\right).$ (32)
@@
$\left(\int\sum_{y}{\bm{Q}}_{t}(x_{t},y)p_{t}(y)\log(p^{\theta}_{t}(x_{t}))dx_{t}+\int\frac{p_{t}(x_{t})}{p^{\theta}_{t}(x_{t})}\sum_{y}{\bm{Q}}_{t}(x_{t},y)p^{\theta}_{t}(y)dx_{t}\right)$ (33)
$\left(\int\sum_{y}{\bm{Q}}_{t}(x_{t},y)\frac{p_{t}(y)}{p_{t}(x_{t})}\log(p^{\theta}_{t}(x_{t}))p(x_{t})dx_{t}+\int p_{t}(x_{t})\sum_{y}{\bm{Q}}_{t}(x_{t},y)\frac{p^{\theta}_{t}(y)}{p^{\theta}_{t}(x_{t})}dx_{t}\right)$ (34)
$\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)\left(\frac{p_{t}(y)}{p_{t}(x_{t})}\log(p^{\theta}_{t}(x_{t}))+\frac{p^{\theta}_{t}(y)}{p^{\theta}_{t}(x_{t})}\right)$ (35)
$\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)\left(\frac{p^{\theta}_{t}(y)}{p^{\theta}_{t}(x_{t})}-\frac{p_{t}(y)}{p_{t}(x_{t})}\log\frac{p^{\theta}_{t}(y)}{p^{\theta}_{t}(x_{t})}\right)$ (36)
$\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)[\frac{p_{t}(y)}{p_{t}(x_{t})}\log p^{\theta}_{t}(y)].$ (37)
@@
$\sum_{x_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)p_{t}(y)\log(p^{\theta}_{t}(y))=\sum_{y}\left(\sum_{x_{t}}{\bm{Q}}_{t}(x_{t},y)\right)p_{t}(y)\log(p^{\theta}_{t}(y))=0,$ (39)
@@
$\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)\left(\frac{p^{\theta}_{t}(y)}{p^{\theta}_{t}(x_{t})}-\frac{p_{t}(y)}{p_{t}(x_{t})}\log\frac{p^{\theta}_{t}(y)}{p^{\theta}_{t}(x_{t})}\right)dt.$ (41)
@@
$H(p_{0})+H(p_{1})=\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)K\left(\frac{p_{t}(y)}{p_{t}(x_{t})}\right)dt,$ (45)
@@
$\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)\left[K\left(\frac{p_{t}(y)}{p_{t}(x_{t})}\right)+\left(s_{\theta}(x_{t},t)_{y}-\frac{p_{t}(y)}{p_{t}(x_{t})}\log{s_{\theta}(x_{t},t)_{y}}\right)\right]dt.$ (47)
@@
$\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y\neq x_{t}}{\bm{Q}}_{t}(x_{t},y)\left[K\left(\frac{p_{t}(y)}{p_{t}(x_{t})}\right)+\left(s_{\theta}(x_{t},t)_{y}-\frac{p_{t}(y)}{p_{t}(x_{t})}\log{s_{\theta}(x_{t},t)_{y}}\right)\right]dt.$ (50)
@@
$\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y\neq x_{t}}{\bm{Q}}_{t}(y,x_{t})dt+H(p_{1},p_{r}),$ (51)
where $\bar{\ell}(a,b)=\left(b-a\log b\right)$.
@@
$\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y}{\bm{Q}}_{t}(x_{t},y)\left(s_{\theta}(x_{t},t)_{y}-\frac{p_{t}(y)}{p_{t}(x_{t})}\log{s_{\theta}(x_{t},t)_{y}}\right)dt+D_{KL}(p_{1}||p_{r}).$ (55)
@@
Proposition 6. In the case of the roulette diffusion with roulette-loglinear noise, $H(p_{r})=0$ and $\int_{0}^{1}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}}\sum_{{\bm{y}}\neq{\bm{x}}_{t}}{\bm{Q}}_{t}({\bm{y}},{\bm{x}}_{t})dt=\left(1-\frac{1-p_{m}}{n-1}\right)\frac{L}{p_{m}}(\epsilon-1)$. For the absorb diffusion, that is when $p_{m}\rightarrow 1$, we have $\int_{0}^{1}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}}\sum_{{\bm{y}}\neq{\bm{x}}_{t}}{\bm{Q}}_{t}({\bm{y}},{\bm{x}}_{t})dt=L(\epsilon-1)$. Finally, for the uniform diffusion, $H(p_{r})=L\log(V)$ and $\int_{0}^{1}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}}\sum_{{\bm{y}}\neq{\bm{x}}_{t}}{\bm{Q}}_{t}({\bm{y}},{\bm{x}}_{t})dt=-\left(1-\frac{1}{V}\right)L\int_{0}^{1}\sigma^{{}^{\prime}}(t)dt$.
@@
$\int_{0}^{1}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}}\sum_{{\bm{y}}\neq{\bm{x}}_{t}}{\bm{Q}}_{t}({\bm{y}},{\bm{x}}_{t})dt=-\mathbb{E}_{t\sim U(0,1)}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\sum_{i=1}^{L}\sum_{{\bm{y}}^{i}\neq{\bm{x}}^{i}_{t}}{\bm{Q}}_{t}^{tok}({\bm{y}}^{i},{\bm{x}}^{i}_{t})$ (58)
$\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\sum_{{\bm{y}}^{i}\neq{\bm{x}}^{i}_{t}}{\bm{Q}}_{t}^{tok}({\bm{y}}^{i},{\bm{x}}^{i}_{t})$ (59)
$\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\sum_{{\bm{x}}^{i}_{t}}p_{t}({\bm{x}}^{i}_{t})\sum_{{\bm{y}}^{i}\neq{\bm{x}}^{i}_{t}}{\bm{Q}}_{t}^{tok}({\bm{y}}^{i},{\bm{x}}^{i}_{t})$ (60)
$\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\sum_{{\bm{x}}^{i}_{0}}p_{0}({\bm{x}}^{i}_{0})\sum_{{\bm{x}}^{i}_{t}}p_{t|0}({\bm{x}}^{i}_{t}|{\bm{x}}^{i}_{0})\sum_{{\bm{y}}^{i}\neq{\bm{x}}^{i}_{t}}{\bm{Q}}_{t}^{tok}({\bm{y}}^{i},{\bm{x}}^{i}_{t})$ (61)
@@
$\int_{0}^{1}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}}\sum_{{\bm{y}}\neq{\bm{x}}_{t}}{\bm{Q}}_{t}({\bm{y}},{\bm{x}}_{t})=$ (62)
$\left(1-\frac{1-p_{m}}{n-1}\right)\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\sum_{{\bm{x}}^{i}_{0}}p_{0}({\bm{x}}^{i}_{0})\sum_{{\bm{x}}^{i}_{t}\neq n}p_{t|0}({\bm{x}}^{i}_{t}|{\bm{x}}^{i}_{0})\sigma_{t}^{{}^{\prime}}=$ (63)
$\left(1-\frac{1-p_{m}}{n-1}\right)\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\sum_{{\bm{x}}^{i}_{0}}p_{0}({\bm{x}}^{i}_{0})\sigma_{t}^{{}^{\prime}}\sum_{{\bm{x}}^{i}_{t}\neq n}p_{t|0}({\bm{x}}^{i}_{t}|{\bm{x}}^{i}_{0})$ (64)
$\left(1-\frac{1-p_{m}}{n-1}\right)\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\sum_{{\bm{x}}^{i}_{0}}p_{0}({\bm{x}}^{i}_{0})\sigma_{t}^{{}^{\prime}}\left(1-p({\bm{x}}^{i}_{t}=n|{\bm{x}}^{i}_{0})\right)$ (65)
$\left(1-\frac{1-p_{m}}{n-1}\right)\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\sum_{{\bm{x}}^{i}_{0}}p_{0}({\bm{x}}^{i}_{0})\sigma_{t}^{{}^{\prime}}e^{-\sigma_{t}p_{m}}=-\left(1-\frac{1-p_{m}}{n-1}\right)L\int_{0}^{1}\sigma_{t}^{{}^{\prime}}e^{-\sigma_{t}p_{m}}dt$ (66)
$\left(1-\frac{1-p_{m}}{n-1}\right)\frac{L}{p_{m}}(e^{-\sigma_{1}p_{m}}-e^{-\sigma_{0}p_{m}})$ (67)
@@
$\mathbb{E}_{t\sim U(0,1)}\sum_{i=1}^{L}\sum_{{\bm{x}}^{i}_{0}}\sum_{{\bm{x}}^{i}_{t}}p({\bm{x}}^{i}_{t},{\bm{x}}^{i}_{0})\sigma^{{}^{\prime}}(t)\left(\frac{1}{V}-1\right)=-L\left(1-\frac{1}{V}\right)\int_{0}^{1}\sigma^{{}^{\prime}}(t)dt.$ (69)